discrete math, please use the K-map to find the simpler circuit with the same output

c) - w ing Fral(x + ) + y+ L+7 十三 4-4-4 74LT (ax+3)+(y+z)
Answer #1

Given ,

Output = Tyz (:1 + 7) + (y + 2)Output = īyz|(1+z)+(+ ) = qyz|(2+y+z) = zīyz + xyyz + xyzz

Output = 0 (TT= yy= zā=0)

Truth Table:

x y z Output
0 0 0 0
0 0 1 0
0 1 0 0
0 1 1 0
1 0 0 0
1 0 1 0
1 1 0 0
1 1 1 0

K Map:

그 00 o ft 10 ||0|0|0 output = o. No gates are requind.The gate gives output always 0. So simplified circuit doesn't require any gates.

Note that since output is always 0, no gates are required. In case if you still want a circuit: (trivial circuit).

3 i/P AND gate: ☆ ON output = 0 O 4 ilp AND: - output o